### v1

# DMol: A Schedule-Driven Diffusion Model for Highly Efficient and Versatile Molecule Generation

Abstract
We introduce a new graph diffusion model for small molecule generation, DMol, which outperforms the state-of-the-art DiGress model in terms of validity by roughly $1.5\%$ across all benchmarking datasets while reducing the number of diffusion steps by at least $10$-fold, and the running time to roughly one half. The performance improvements are a result of a careful change in the objective function and a “graph noise” scheduling approach which, at each diffusion step, allows one to only change a subset of nodes of varying size in the molecule graph. Another relevant property of the method is that it can be easily combined with junction-tree-like graph representations that arise by compressing a collection of relevant ring structures into supernodes. Unlike classical junction-tree techniques that involve VAEs and require complicated reconstruction steps, compressed DMol directly performs graph diffusion on a graph that compresses only a carefully selected set of frequent carbon rings into supernodes, which results in straightforward sample generation. This compressed DMol method offers additional validity improvements over generic DMol of roughly $2\%$, increases the novelty of the method, and further improves the running time due to reductions in the graph size.
Machine Learning, ICML
1 Introduction
Graphs are fundamental data structures used to model a wide range of complex interactions, including molecules and drugs, biological pathways, and social and co-purchase networks (Wu et al., 2021; Zhang et al., 2024; Yang et al., 2024; Waikhom & Patgiri, 2021). Therefore, the ability to generate graphs that accurately capture domain-specific distributions is of great significance (Tsai et al., 2023; You et al., 2018; Gamage et al., 2020). While generative models for images and texts are already being used with great success (Dhariwal & Nichol, 2021; Ho et al., 2022; Waikhom & Patgiri, 2021), graph generation remains a challenging frontier due to the discrete nature of the models, the need for permutation invariance and other distinctive properties (Velikonivtsev et al., 2024; Wang et al., 2023). This is particularly the case when trying to generate graph samples that mimic and innovate molecular compounds (Jin et al., 2018; Simonovsky & Komodakis, 2018; Mitton et al., 2021; Vignac & Frossard, 2021; Gómez-Bombarelli et al., 2018; Kwon et al., 2020; Wu et al., 2024; You et al., 2018; Liao et al., 2019; Segler et al., 2018; Popova et al., 2019; Madhawa et al., 2019; Zang & Wang, 2020; Luo et al., 2021b), in which case the generated samples can fail to obey necessary biochemical and physical constraints.
One way to resolve some of these challenges is to use diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Fan et al., 2023; Chen et al., 2024; Trippe et al., 2023; Haefeli et al., 2023; Niu et al., 2020b; Huang et al., 2022; Liu et al., 2024c) which rely on a gradual noise injection process that perturbs the input samples and then reverses the process using a learned denoising network. In the context of small molecule drug generation, existing methods focus on discrete diffusion in order to avoid the loss of key structural properties that arise when embedding graph structures into continuous spaces (Niu et al., 2020a; Jo et al., 2022b). Discrete diffusion models incorporate “transition mechanisms” that explicitly consider the categorical attributes of nodes and edges, ensuring structural fine-graining and integrity during the noise injection process (Liu et al., 2024b; Xu et al., 2024). They also allow for continuous and discrete denoising steps that diversify the outputs (Huang et al., 2022; Vignac et al., 2023; Jo et al., 2022a). However, they also typically require a very large number of diffusion steps, resulting in large running times and extremely slow graph generation (Kong et al., 2023). More importantly, they lead to structures that have low chemical validity even when measured through highly coarse SMILES nomenclature constraints (Hoogeboom et al., 2022; Liu et al., 2021).
The state-of-the-art graph diffusion method for generating small molecules, DiGress (Vignac et al., 2023), relies on a “cosine” noise injection mechanism that controls the amount of perturbations during the different diffusion steps. In the process, every node and edge in the graph is allowed to be modified at each step. DiGress evaluates the generated drug samples through validity, novelty, and uniqueness scores that provide a limited measure of the utility of molecules. Unfortunately, its running time is excessive and almost all samples generated in a run have unacceptably low biochemical scores seldom tested in the literature. Furthermore, in its current form, and unlike some non-diffusion based methods (Jin et al., 2018; Simonovsky & Komodakis, 2018; Mitton et al., 2021; Vignac & Frossard, 2021; Gómez-Bombarelli et al., 2018; Kwon et al., 2020; Wu et al., 2024; You et al., 2018; Liao et al., 2019; Segler et al., 2018; Popova et al., 2019), DiGress cannot preserve the structures and frequencies of important network motifs (e.g., carbon rings, which are the most important building blocks of molecules).
We describe a new small molecule drug diffusion model, (Compressed) DMol, which significantly outperforms DiGress across all benchmarking datasets in terms of running time, SMILES validity, and novelty scores and at the same time, allows for network motif preservation via compressed representations of subsets of rings. Its main features are:
1. DMol directly relates the diffusion time-steps with the number of nodes and edges that are allowed to undergo state changes in that step. This approach effectively increases the learning rate of the model as it learns increasingly larger submodels in a gradual manner. It also significantly reduces the time required for molecule generation. Furthermore, it uses a new objective function that penalizes mismatches in the count of nodes and edges during forward and backward steps. Unlike DiGress, DMol also exhibits a more flexible control of the noise schedule/distribution.
2. DMol can operate on graphs which “compress” important node motifs into supernodes. In compressed DMol, the key idea is to combine diffusion models with compressed motif representations of molecules akin to the ones used in JT-VAE (Jin et al., 2018). But unlike JT-VAE which requires sophisticated and time-consuming decoding, compressed DMol only uses a small number of carbon ring structures as super nodes and performs diffusion on the compacted graphs directly. Since the rings are chosen in such a way that only one type of bond is possible during reconstruction, the decoding is extremely simple. Ring compression differs from ring freezing as it still allows one ring structure to be changed to another. The validity improvement of compressed DMol over DiGress is $>3\%$ over all benchmarking datasets while the number of diffusion steps remains an order of magnitude smaller.
3. DMol also uniformly improves the performance of DiGress under previously unexplored validity metrics, such as ChEMBL likeness (which indicates how likely the generated molecule is to have similar biological activity as an already known drug in the ChEMBL molecule repository) and QED (quantitative estimate of drug-likeness), which is integrated into RdKit and measures chemical “desirability” (Landrum et al., 2006).
2 Related Works
Viewing small molecules as graphs has inspired a myriad of generative models for drug design, including JT-VAE (Jin et al., 2018), MoFlow (Zang & Wang, 2020), CDGS (Huang et al., 2023), EDM (Hoogeboom et al., 2022), etc. (Cornet et al., 2024; Shi et al., 2021; Xu et al., 2024; Chen et al., 2023; Wu et al., 2022; Bao et al., 2023). Furthermore, numerous surveys on the subject are readily available (Tang et al., 2024; Du et al., 2022; Yang et al., 2024). We partition and review prior works on generative models for drug discovery based on the methodology used.
2.1 Traditional Methods
Traditional approaches to graph-based molecular generation relied on the use of Variational Autoencoders (VAEs) as generative models (Jin et al., 2018; Simonovsky & Komodakis, 2018; Mitton et al., 2021; Vignac & Frossard, 2021; Gómez-Bombarelli et al., 2018; Kwon et al., 2020; Wu et al., 2024). VAEs comprise an encoder that maps the input graph into a latent space and a decoder that reconstructs the graph from the latent embedding. Many methods also involve autoregressive models, which generate molecular graphs by sequentially predicting the next element in the sequence based on previous outputs (You et al., 2018; Liao et al., 2019; Segler et al., 2018; Popova et al., 2019). On the other hand, flow-based methods leverage the concept of normalizing flows for graph generation (Madhawa et al., 2019; Zang & Wang, 2020; Luo et al., 2021b; Lippe & Gavves, 2021; Shi* et al., 2020). These techniques apply a series of invertible transformations to model complicated distributions based on simpler ones (such as a Gaussian). Additionally, Generative Adversarial Networks (GANs) approaches have also been used for molecule generation (De Cao & Kipf, 2018; Łukasz Maziarka et al., 2019; Martinkus et al., 2022a). GANs include a generator, which learns to create realistic molecular graphs, and a discriminator, which learns to differentiate between real and generated samples.
2.2 Diffusion-Based Methods
Diffusion models for graph generation can be broadly categorized as follows. Denoising Diffusion Probabilistic Models (DDPM) utilize a Markov chain for the diffusion process (Vignac et al., 2023; Xu et al., 2022; Liu et al., 2024b; Xu et al., 2024; Liu et al., 2024a; Hoogeboom et al., 2022; Jo et al., 2024). Score-Based Generative Models (SGM) leverage score matching techniques to model the data distribution, and generate graphs by reversing a diffusion process guided by the score of the data distribution (Luo et al., 2021a; Chen et al., 2023; Wu et al., 2022). Furthermore, by replacing discrete time steps with continuous time, one can use stochastic differential equations to model the diffusion and denoising processes (Jo et al., 2022a; Lee et al., 2023; Huang et al., 2023; Bao et al., 2023). Furthermore, studies have shown that different noise schedules may significantly affect molecule generation quality (Shi et al., 2025; Nichol & Dhariwal, 2021). Comprehensive surveys on the subjects include (Mengchun Zhang et al., 2023; Fan et al., 2023; Chen et al., 2024).
DiGress (Vignac et al., 2023) uses a denoising diffusion probabilistic model within a discrete state space and is considered one of the state-of-the-art models for molecular generation tasks. It represents graphs in the discrete space of node and edge attributes by assigning categorical labels to each node and edge. DiGress models the noise addition process as a Markov process, where each node and edge label evolves independently of the others. This assumption mirrors the approach in image-based diffusion models. Denoising is performed using a graph transformer network, trained by minimizing the cross-entropy loss between the predicted probabilities for nodes and edges and the ground-truth graph. The transformer network reconstructs a denoised graph from a noisy input. To generate new graphs, a noisy graph is first sampled according to the model’s limiting distribution, and the trained denoising transformer is then used to produce a graph with the desired properties. The DiGress model is discussed further in Section 5.
3 DMol Diffusion
3.1 Review of the Graph Diffusion Model
The focal concept in diffusion models is the noise model $q$, designed to enable $T$ forward diffusion steps. Graph diffusion relies on progressively introducing noisy perturbations to an undirected graph $\mathcal{G}^{0}$, encoded as $\textbf{z}^{0}$, resulting in a sequence of increasingly noisy data points $(\textbf{z}^{1},\dots,\textbf{z}^{T})$. The noise addition process is Markovian, which translates to $q(\textbf{z}^{1},\dots,\textbf{z}^{T}\mid\textbf{z}^{0})=\prod_{t=1}^{T}q(\textbf{z}^{t}\mid\textbf{z}^{t-1})$.
More precisely, the encoding of a graph involves categorical information about the nodes $V$ and vertices $E$ of the graph $\mathcal{G}=(V,E).$ The assumptions are that each node belongs to one of $a$ classes and is represented as a one-hot vector in $\mathbb{R}^{a}$. The same is true for edges, which are assumed to belong to one of $b$ classes (with one class indicating the absence of the edge). We follow the notation from (Vignac et al., 2023), where $\textbf{x}_{i}$ is used to denote the one-hot encoding of the class of node $i$. The one-hot encodings are arranged in a node matrix $\bf{X}$ of dimensions $n\times a$. Again, in a similar manner, edges are encoded into a tensor $\bf{E}$ of dimension $n\times n\times b$. The Markovian state space for each node and edge is the set of classes of the nodes and edges, respectively. For simplicity, and with a slight abuse of notation, $\textbf{x}_{l}$ is used to denote the class (state) of node $l$, while $\textbf{e}_{ls}$ is used to denote the class (state) of edge $ls$.
Diffusion is performed separately on nodes and edges, while the noisy perturbations in the forward process are governed by a state transition matrix. For node-level noise we use transition matrices $(\textbf{Q}_{X}^{1},\dots,\textbf{Q}_{X}^{T})$, where each entry $[\textbf{Q}_{X}^{t}]_{ij}$ defines the probability of an arbitrary node transitioning from state $i$ to state $j$ at time step $t$. A similar model is used for edges, where the transition matrices $(\textbf{Q}_{E}^{1},\dots,\textbf{Q}_{E}^{T})$ describe the probabilities $[\textbf{Q}_{E}^{t}]_{ij}$ of an arbitrary edge transitioning from state $i$ to state $j$. Formally, $[\textbf{Q}_{X}^{t}]_{ij}=q(x^{t}=j\mid x^{t-1}=i)$ and $[\textbf{Q}_{E}^{t}]_{ij}=q(e^{t}=j\mid e^{t-1}=i)$, where $x$ and $e$ correspond to a generic node and edge, while the superscript indicates the time stamp.
Consequently, the transition probability at time $t$ is given by $q(\textbf{z}^{t}\mid\textbf{z}^{t-1})={\textbf{z}^{t-1}}\textbf{Q}^{t}$, where z stands for either x or e and the subscript of Q is set accordingly. Adding noise to a graph $\mathcal{G}^{t-1}$ at time $t-1$ results in a new graph $\mathcal{G}^{t}=(\textbf{X}^{t},\textbf{E}^{t}),$ and involves sampling each node and edge class from a categorical distribution succinctly described as $q(\mathcal{G}^{t}\mid\mathcal{G}^{t-1})=\big{(}\textbf{X}^{t-1}\textbf{Q}_{X}^{t},\textbf{E}^{t-1}\textbf{Q}_{E}^{t}\big{)}$ and $q(\mathcal{G}^{t}\mid\mathcal{G}^{0})=\big{(}\textbf{X}^{0}\overline{\textbf{Q}}_{X}^{t},\textbf{E}^{0}\overline{\textbf{Q}}_{E}^{t}\big{)}$, where $\overline{\textbf{Q}}_{X}^{t}=\textbf{Q}_{X}^{1}\dots\textbf{Q}_{X}^{t}$ and $\overline{\textbf{Q}}_{E}^{t}=\textbf{Q}_{E}^{1}\dots\textbf{Q}_{E}^{t}$. The state transition matrices are invertible, which ensures the reversibility of the noise model: the graphs at time $t$ and time 00 can be transformed into each other using the state transition matrix. Reversibility guarantees consistency between the forward and backward processes and simplifies denoising in the backward process.
A common issue with most known diffusion models is the large number of steps needed for convergence (Hang et al., 2023; Wang et al., 2024). Although there are typically no strict time constraints for molecule generation, for large training sets the computational complexity may be overwhelming (see also the Results section). Hence, one important question is how to reduce the number of diffusion steps for graph generation, and at the same time, improve the utility of the model. These issues are addressed via our new DMol diffusion model.
3.2 DMol Forward Noise Adding Strategy
The key idea behind our approach is as follows: at each forward diffusion step $t$, we randomly select $N(t)$ nodes and $M(t)$ edges from $\mathcal{G}^{0}$ contained in the complete subgraph induced by the selected edges (note that the absence of an edge is represented by a special label so that we are effectively dealing with a complete graph), and change their classes according to fixed state transition matrices $\textbf{Q}_{X}$ and $\textbf{Q}_{E}$, respectively, to obtain the graph $\mathcal{G}^{t}$. Here, $N(t)$ and $M(t)$ are functions of the time step index $t$. Note that we effectively run the diffusion process on randomly selected subsets of nodes and edges, which is one of the key differences when compared to DiGress.
Figure 1: The forward noise addition process of DMol. At each time step $t$, we randomly selected a predetermined number of nodes and added noise only to the class labels of the nodes and a fixed number of edges in the subgraph induced by the nodes. Both selection numbers are controlled by the time index $t$. The top row depicts how the number of selected nodes $N(t)$ and edges $M(t)$ of the selected subgraph evolves with $t$ (we set $n=6,k=2,r=0.2$ and the values need to be floored to obtain integers).
Figure 2: DMol diffusion relies on randomly selecting $N(t)$ nodes and $M(t)$ edges at each time step $t$ and diffusing their labels to generate a noisy graph $\mathcal{G}^{t}$. The denoising network $\phi_{\theta}$ learns to predict the denoised graph from $\mathcal{G}^{t}$. During inference, the denoised graph is combined with $q(\mathcal{\hat{G}}^{t-1}|\mathcal{\hat{G}}^{t},\mathcal{\hat{G}}^{0})$ to predict $\mathcal{\hat{G}}^{t-1}$.
Algorithm 1 Forward Process
  Input: Graph $\mathcal{G}^{0}=(\textbf{X}^{0},\textbf{E}^{0})$, state transition matrices $\textbf{Q}_{X}\&\textbf{Q}_{E}$, node&edge selection functions $N(\cdot)\&M(\cdot)$.
  Sample $t\in\{1\dots T\}$.
  Sample $\mathcal{\tilde{G}}^{t}=(\tilde{\textbf{X}}^{t},\tilde{\textbf{E}}^{t})$ from $\textbf{X}^{t}\textbf{Q}_{X}\times\textbf{E}^{t}\textbf{Q}_{E}$.
  Calculate the number of nodes and edges, $u=N(t)$ and $v=M(t)$, on which to diffuse.
  Randomly select $u$ nodes and $v$ edges from the subgraph induced by the $u$ nodes and generate “filter” matrices $\textbf{K}_{x}$ and $\textbf{K}_{e}$ which serve as indicators of the selections.
  Set $\mathcal{G}^{t}=\mathcal{G}^{0}$ and replace $\textbf{K}_{x}\textbf{X}^{t}$ with $\textbf{K}_{x}\tilde{\textbf{X}}^{t}$, and $\textbf{K}_{e}\textbf{E}^{t}$ with $\textbf{K}_{e}\tilde{\textbf{E}}^{t}$.
[…]
Additionally, our forward process constrains the class changes of edges to the subgraph induced by the nodes that underwent state change. This ensures that the sets of altered nodes and edges are codependent. Still, once the selections are made, the state changes of nodes and edges are governed by independent processes.
Next, we turn our attention to the choice of transition probability matrices. The results of DiGress (Vignac et al., 2023) have shown that making the transition probability from state $i$ to state $j$ proportional to the marginal probability of class $j$ within the training dataset results in more effective learning of the true data distribution compared to when using uniform transitions. We follow the same procedure and define the marginal probabilities of node classes in the training dataset as $\mathbf{m}_{x}\in\mathbb{R}^{1\times a}$ and those of edge classes as $\mathbf{m}_{e}\in\mathbb{R}^{1\times b}$, respectively. Hence,
$\textbf{Q}_{X}[i,j]=\begin{cases}\frac{\mathbf{m}_{x}[j]}{\sum_{\ell\neq i}\mathbf{m}_{x}[\ell]},&\text{if }i\neq j,\\ 0,&\text{if }i=j;\end{cases}$
$\textbf{Q}_{E}[i,j]=\begin{cases}\frac{\mathbf{m}_{e}[j]}{\sum_{\ell\neq i}\mathbf{m}_{e}[\ell]},&\text{if }i\neq j,\\ 0,&\text{if }i=j.\end{cases}$
During the forward process, we compute the distributions of all node and edge classes after state transitions. However, we only focus on the distributions corresponding to the selected $N(t)$ nodes and $M(t)$ edges, sampling from these distributions to update their states. We preserve the states of the remaining nodes from the previous step. Algorithm 1 and Figure 2 illustrate our forward process.
3.3 Sampling Process
During the sampling process, we first sample a random graph based on the marginal distributions of nodes and edges at $t=T$. This random graph is then used as the input to the denoising network $\phi_{\theta}$ to predict a denoised graph. Based on the prediction, we use the noise model to obtain the noisy graph at $t=T-1$. Subsequently, this noisy graph is fed into the denoising network, and the above process is repeated iteratively until a new graph is sampled.
To start the sampling process, we need to compute the marginal distributions of nodes and edges at $t=T$. At time $t=T$, a node class $i$ at $t=T$ is a result of transitions from some other node classes at $t=0$. As a result, the marginal probability of the $i$-th node class at time $t=T$ equals
$p^{x,T}_{i}=p^{x}_{i}\sum_{j}\frac{p^{x}_{j}}{1-p^{x}_{j}},$
where $p^{x}_{i}$ and $p^{x}_{j}$ are the marginal probabilities of node classes $i$ and $j$ at time $t=0$, respectively. For edges, we similarly have that the marginal probabilities of the $i$-th edge class at $t=T$ equals
$p^{e,T}_{i}=(1-r)\,p^{e}_{i}+r\,p^{e}_{i}\sum_{l}\frac{p^{e}_{l}}{1-p^{e}_{l}},$
where $p^{e}_{i}$ and $p^{e}_{l}$ are the marginal probabilities of edge classes $i$ and $l$ at $t=0$.
Algorithm 2 Sampling Process
  Input: Denoising Network $\phi_{\theta}$, node class distribution $\mathbf{m}_{x}^{T}=[p_{i}^{x,T}]$, edge class distribution $\mathbf{m}_{e}^{T}=[p_{k}^{e,T}]$, state transition matrices $\textbf{Q}_{X}\&\textbf{Q}_{E}$.
  Sample $\mathcal{\hat{G}}^{T}$ from $\mathbf{m}_{x}^{T}\times\mathbf{m}_{e}^{T}$.
  for $t=T$ to $1$ do
     $z\leftarrow f(\mathcal{\hat{G}}^{t},t)$DiGress structural and spectral features, as explained in (Vignac et al., 2023) Appendix B.2}.
     $\hat{p}^{X},\hat{p}^{E}\leftarrow\phi_{\theta}(\mathcal{\hat{G}}^{t},z,t,R_{x},R_{e})$Prediction}.
     Sample $\mathcal{\hat{G}}^{0}=(\hat{\textbf{X}}^{0},\hat{\textbf{E}}^{0})$ from $\hat{p}^{X}\times\hat{p}^{E}$.
     Sample $\tilde{\mathcal{G}}^{t-1}=(\tilde{\textbf{X}}^{t-1},\tilde{\textbf{E}}^{t-1})$ from $\hat{\textbf{X}}^{0}\textbf{Q}_{X}\times\hat{\textbf{E}}^{0}\textbf{Q}_{E}$.
     Calculate $u=N(t-1)$, $v=M(t-1)$.
     Randomly select $u$ nodes and $v$ edges and generate corresponding “filter” matrices $\textbf{K}_{x}$ and $\textbf{K}_{e}$.
     Set $\mathcal{\hat{G}}^{t-1}=\mathcal{\hat{G}}^{0}$ and replace $\textbf{K}_{x}\hat{\textbf{X}}^{t-1}$ with $\textbf{K}_{x}\tilde{\textbf{X}}^{t-1}$, $\textbf{K}_{e}\hat{\textbf{E}}^{t-1}$ with $\textbf{K}_{e}\tilde{\textbf{E}}^{t-1}$.
  end for
After sampling the random graph, we use the denoising model to predict the clean graph and continue this process iteratively. At time step $t$, the input to the denoising model includes not only the noisy graph and the time $t$ but also two additional functions: $R_{x}(t)=N(t)/n$ and $R_{e}(t)=M(t)/(0.5\,n\,(n-1))$, which describe the proportion of nodes and edges in the noisy graph at time $t$ that undergo state transitions relative to the total number of nodes and edges, respectively. They indicate to the denoising network that the input noisy graph and the predicted graph should differ by $N(t)$ nodes and $M(t)$ edges, thereby helping the network make more accurate predictions. The sampling steps are listed in Algorithm  2. Figure  8 in the Appendix is an example of the sampling process from a random graph to a denoised graph.
The loss we use to train the denoising network differs from that of DiGress in so far that it contains additional penalties,
$l(\hat{\textbf{p}}^{\mathcal{G}},\mathcal{G})$ $=\sum_{1\leq i\leq n}\text{CE}(\textbf{x}_{i},\hat{\textbf{p}}^{X}_{i})+\lambda_{1}\sum_{1\leq i,j\leq n}\text{CE}(\textbf{e}_{ij},\hat{\textbf{p}}^{E}_{ij})$
$+\lambda_{2}\,\text{MSE}\big{(}D(\text{argmax}(\hat{\textbf{p}}^{X});\text{argmax}(\textbf{X})),N(t)\big{)}$
$+\lambda_{3}\,\text{MSE}\big{(}D(\text{argmax}(\hat{\textbf{p}}^{E});\text{argmax}(\textbf{E})),M(t)\big{)}.$
[…]
We also briefly comment on the equivariance property of the diffusion model and the invariance of the loss function used by DMol.
Proposition 3.1 (Equivariance).
The DMol diffusion model is permutation equivariant.
The result easily follows since our denoising model used the same graph transformer as DiGress, which is permutation equivariant.
Proposition 3.2 (Invariance of the Loss Function).
The DMol loss function is permutation invariant.
Our loss function comprises two components, a cross-entropy loss and an MSE loss.
For the cross-entropy loss terms,
$l_{CE}(\hat{\textbf{p}}^{\mathcal{G}},\mathcal{G})=\sum_{1\leq i\leq n}\text{CE}(\textbf{x}_{i},\hat{\textbf{p}}^{X}_{i})+\lambda_{1}\sum_{1\leq i,j\leq n}\text{CE}(\textbf{e}_{ij},\hat{\textbf{p}}^{E}_{ij}),$
Since the overall cross-entropy loss is calculated by summing up the individual cross-entropy values for each node and edge, and the sum operation is permutation invariant, then the cross-entropy loss is also invariant.
We arrive at a similar conclusion for the MSE terms in $l_{MSE}(\hat{\textbf{p}}^{\mathcal{G}},\mathcal{G})$,
$\lambda_{2}\,\text{MSE}\big{(}D(\text{argmax}(\hat{\textbf{p}}^{X});\text{argmax}(\textbf{X})),N(t)\big{)}$
$+$ $\lambda_{3}\,\text{MSE}\big{(}D(\text{argmax}(\hat{\textbf{p}}^{E});\text{argmax}(\textbf{E})),M(t)\big{)}$
where invariance follows based on the definition of $D(\cdot;\cdot)$ and the fact that the loss is computed separately for nodes/edges with different indices and then summed.
4 Ring Compression and Compressed DMol
Inspired by JT-VAE (Jin et al., 2019), we incorporate into DMol a ring compression process, depicted in Figure 3. The process is designed to better preserve motifs in molecular graphs, which are predominantly carbon rings. Specifically, we convert a small (roughly $10$) predefined number of most frequently occurring carbon rings that can only form single bonds through their available carbons into supernodes, introducing in the process a new node class to represent the compressed rings. The compressed graph representations are directly fed into our DMol diffusion model. During the sampling process, the supernodes are converted back to their respective carbon rings.
Figure 3: In molecule graphs, frequently occurring ring structures that also allow for only single bonds with carbon atoms are compressed into supernodes during preprocessing. Each supernode comes with its own label. During DMol diffusion, supernodes are either converted into other classes of supernodes (e.g., rings can be converted into other classes of rings) or into “simple” nodes, (e.g., individual atoms), and vice versa. During the sampling process, the supernode is decoded back into its corresponding ring, and integrated into the graph by using the only possible bond structure. Since molecular graphs are sparse, it is highly unlikely to exceed two bonds per supernode.
There are significant differences between our node compression approach and that of junction tree encoding in JT-VAE (Jin et al., 2019). First, the JT approach requires finding a minimum spanning tree during the encoding process. Since VAEs do not directly generate graphs, one has to perform complicated decoding. For example, each compressed molecular substructure can form many different types of bonds with different atoms in the reconstruction (i.e., reconstruction is not unique). The options are ranked according to a special scoring function and the overall process is computationally demanding. In contrast, our method simplifies the encoding and decoding processes by directly transforming only the most common carbon rings into supernodes which also allow only one or two bonding options. It is important to note that ring compression differs from node freezing since we allow different classes of rings to be converted into each other or simple atoms. The ring structures used in Compressed DMol for different datasets are available in the Appendix, Figures 456.
5 DiGress vs DMol
We next compare DiGress and DMol. First, we observe that DiGress also constructs state transition matrices based on marginal distributions, with the key difference with respect to DMol being that its transition matrices vary over time as
$\textbf{Q}_{X}^{t}=\alpha^{t}\mathbf{I}+\beta^{t}\mathbf{1}_{a}\mathbf{m}_{x},\quad\textbf{Q}_{E}^{t}=\alpha^{t}\mathbf{I}+\beta^{t}\mathbf{1}_{b}\mathbf{m}_{e},$
where $\alpha^{t}$ and $\beta^{t}$ are time-dependent scaling parameters, while $\textbf{Q}_{X}^{t}$ and $\textbf{Q}_{E}^{t}$ are the state transition matrices used to transition from $\mathcal{G}^{t-1}$ to $\mathcal{G}^{t}$. However, since the forward noise adding process is Markovian, DiGress directly utilizes $\overline{\textbf{Q}}_{X}^{t}$ and $\overline{\textbf{Q}}_{E}^{t}$ to achieve the transition from $\mathcal{G}^{0}$ to $\mathcal{G}^{t}$, where $\overline{\textbf{Q}}_{X}^{t}=\textbf{Q}_{X}^{1}\dots\textbf{Q}_{X}^{t}$ and $\overline{\textbf{Q}}_{E}^{t}=\textbf{Q}_{E}^{1}\dots\textbf{Q}_{E}^{t}$.
Since $(\mathbf{1}\mathbf{m})^{2}=\mathbf{1}\mathbf{m}$, we have
$\overline{\textbf{Q}}_{X}^{t}=\overline{\alpha}^{t}\mathbf{I}+\overline{\beta}^{t}\mathbf{1}_{a}\mathbf{m}_{x},\overline{\textbf{Q}}_{E}^{t}=\overline{\alpha}^{t}\mathbf{I}+\overline{\beta}^{t}\mathbf{1}_{b}\mathbf{m}_{e},$
where $\overline{\alpha}^{t}=\prod_{\tau=1}^{t}\alpha^{\tau}$ and $\overline{\beta}^{t}=1-\overline{\alpha}^{t}$. Additionally, DiGress also adopts a cosine schedule for $\overline{\alpha}^{t}$$\overline{\alpha}^{t}=\cos^{2}(0.5\pi(t/T+c)/(1+c)),$ where $c>0$ is a constant.
In DMol, for a graph dataset with $n$ nodes, the number of nodes undergoing state changes at time step $t$ equals $N(t)=(1-\alpha)\,n$. At the same time step in Digress, the expected number of nodes undergoing state changes equals $\sum_{i}np^{x}_{i}(1-\alpha)(1-p^{x}_{i})$, where $p^{x}_{i}$ represents the marginal probability of node class $i$. Therefore, at the same time step, the number of nodes modified by DMol is $\frac{1}{\sum_{i}p^{x}_{i}(1-p^{x}_{i})}$ times that expected from Digress (the derivations can be found in Appendix C.2). Moreover, among the nodes whose states are changed by DMol, the proportion of nodes with class $i$ is $p^{x}_{i}$, whereas in DiGress, this proportion equals $p^{x}_{i}(1-p^{x}_{i})$. The ratio of these fractions equals $1-p^{x}_{i}$.
For edges, a similar analysis can be applied, leading to the conclusion that at the last time step $T$, the ratio of the number of edges modified by our noise model and that modified by Digress equals $\frac{r}{\sum_{j}p^{e}_{j}(1-p^{e}_{j})}$, where $r$ is the hyperparameter mentioned in Section 3.2. At each time step, this ratio needs to be adjusted by a correction factor that accounts for the fact that our method only selects edges from the complete graph induced by the chosen nodes. This correction factor is the ratio of the number of edges in the induced subgraph and the total number of edges. Additionally, among the modified edges, the proportion of edges with class index $j$ differs from the same in DiGress by a constant factor of $1-p^{e}_{j}$, where $p^{e}_{j}$ equals the marginal probability of edge class $j$. As a result, DMol can be made to replicate the average “behaviour” of DiGress by simply adjusting specific hyperparameters, whereas the opposite claim is not true. Comparisons of the noise evolution distribution, efficiency and expressivity of DMol and Digress are available in Appendix B and Appendix C.
6 Experiments
We present next the results of running DMol and Compressed DMol [^1] The link to the code: https://github.com/liekon/Discrete-Graph-Generation.git on several benchmarking datasets. For experiments pertaining to other graph models (e.g., SBMs and planar graphs), the reader is referred to Appendix 6.
Benchmarks. We compared DMol and compressed DMol to several graph generation methods. The benchmarking models used in the experiments include GraphRNN (You et al., 2018), GRAN (Liao et al., 2019), GraphVAE (Simonovsky & Komodakis, 2018), GT-VAE (Mitton et al., 2021), Set2GraphVAE (Vignac & Frossard, 2021), SPECTRE (Martinkus et al., 2022b), GraphNVP (Madhawa et al., 2019), GDSS (Jo et al., 2022b), VAE (Gómez-Bombarelli et al., 2018), JT-VAE (Jin et al., 2019), GraphINVENT (Mercado et al., 2021), LSTM (Segler et al., 2018), NAGVAE (Kwon et al., 2020), MCTS (Jensen, 2019) and DiGress (Vignac et al., 2023).
Data. We tested the performance of different generative models on the QM9 (Wu et al., 2018), MOSES (Polykovskiy et al., 2020), and GUACAMOL (Brown et al., 2019) datasets. QM9 is a small dataset that consists of molecules containing up to $9$ heavy atoms; we used $100$K molecules for training, $20$K for validation, and $13$K for testing. Both MOSES and GuacaMol contain millions of molecules, with the number of heavy atoms inside a single molecule bounded by $26$ and $88$, respectively. We use 85% of the molecules for training, 5% for validation, and 10% for testing.
Setup. For each molecule dataset, we run DMol and compressed DMol. For the compressed algorithm, we choose a different (relatively small) number of ring types to be converted into supernodes. The selected ring structures can be found in the Appendix A. We set the hyperparameters of DMol in both setting to $k=2$, $r=0.2$, $\lambda_{1}=5$, $\lambda_{2}=1$, $\lambda_{3}=1$. The number of training epochs is set to $500$. The experiments were conducted using NVIDIA H100 GPUs. The QM9 runs were executed on a single GPU, while the MOSES and GUACAMOL experiments were processed in parallel using $4$ GPUs.
Table 1: Performance comparison on the QM9 dataset.
Model V↑ V.U.↑ V.U.N.↑
GraphVAE 56.1 42.8 26.5
GT-VAE 74.8 16.5 15.6
Set2GraphVAE 60.0 56.8 -
SPECTRE 87.5 31.4 29.0
GraphNVP 83.5 18.4 -
GDSS 96.0 94.6 -
DiGress 97.8 95.2 31.8
CDiGress(Ours) 98.0 95.9 52.8
DMol(Ours) 98.2 95.8 68.6
CDMol(Ours) 98.3 96.0 75.1
Table 2: Performance comparison on MOSES. All graph-based methods except DiGress and DMol have hard-coded rules to ensure high validity. Note that we report slightly lower V values for DiGress compared to the original results, as obtained from our own experiments.
Model V↑ U↑ N↑ Filters↑ FCD↓ SNN↑ Scaf↑
VAE 97.2 99.4 68.6 98.9 0.56 0.62 6.1
JT-VAE 100 100 99.8 96.9 1.08 0.52 9.8
GraphINVENT 96.1 99.5 - 95.8 1.20 0.56 12.4
DiGress 84.8 100 94.5 97.2 1.18 0.55 14.6
CDiGress (Ours) 85.8 100 96.2 97.0 1.16 0.56 14.9
DMol (Ours) 85.6 100 100 97.4 1.15 0.55 14.9
CDMol (Ours) 87.8 100 100 97.8 1.12 0.58 14.8
Table 3: Performance comparison on GUACAMOL.
Model V↑ U↑ N↑ KL div↑ FCD↑
LSTM 96.1 100 91.0 99.3 91.5
NAGVAE 93.2 95.4 100 39.5 1.0
MCTS 100 100 95.2 82.9 1.8
DiGress 84.5 100 99.8 93.1 68.4
CDiGress (Ours) 85.6 100 100 93.0 68.8
DMol (Ours) 85.4 100 100 93.5 68.6
CDMol (Ours) 86.7 100 100 94.2 69.3
Table 4: Performance comparison of molecular metrics on MOSES and GUACAMOL. CL stands for CLscores, ST stands for sample time, and ATRT stands for approximate training time.
Dataset Model CL↑ QED↑ ST(s)↓ ATRT(days)↓
Moses DiGress 4.4965 0.8024 68.24 5.5
CDiGress (Ours) 4.4984 0.8036 75.46 5.5
DMol (Ours) 4.5002 0.8069 5.28 3.0
CDMol (Ours) 4.5033 0.8055 1.52 3.0
Guacamol DiGress 4.0412 0.5650 70.82 6.0
CDiGress (Ours) 4.1902 0.5698 81.36 6.0
DMol (Ours) 4.0425 0.5728 5.49 4.0
CDMol (Ours) 4.2231 0.5786 1.84 4.0
Evaluation Metrics. The metrics used in the experiments include validity, uniqueness, and novelty. Validity (V) measures the proportion of generated molecules that pass basic valency checks (e.g., correspond to valid SMILES files). Uniqueness (U) quantifies the proportion of generated molecules that are distinct. Novelty (N) represents the percentage of generated molecules that do not appear in the training dataset. Furthermore, V.U. stands for the product of validity and uniqueness scores, while V.U.N. stands for the product of validity, uniqueness, and novelty scores.
Since MOSES and GUACAMOL are benchmarking datasets, they introduce their own set of metrics for reporting results. In addition to V, U, and N, MOSES incorporates a filter score, the Frechet ChemNet Distance (FCD) (Preuer et al., 2018), SNN, and scaffold similarity scores, while GUACAMOL uses FCD and KL divergence. These metrics are defined as follows: The filter score measures the proportion of molecules that pass the same filters used to construct the test set. The FCD score measures the similarity between molecules in the training and test sets using embeddings learned by a neural network. SNN quantifies the similarity of a molecule to its nearest neighbor based on the Tanimoto distance. Scaffold similarity compares the frequency distributions of so-called Bemis-Murcko scaffolds. KL divergence measures the differences in the distributions of various physicochemical descriptors.
The metrics mentioned above still have certain limitations. Since these metrics assess the overall quality of the generation process, but do not provide insights into the performance of individual molecules. To address this limitation, we also use the ChEMBL-Likeness Score (CLscore) (Bühlmann & Reymond, 2020) and Quantitative Estimation of Drug-likeness (QED) (Bickerton et al., 2012) score. The CLscore is calculated by considering how many substructures in a molecule also occur in the drug-like dataset ChEMBL (Gaulton et al., 2016). QED evaluates drug-likeness by considering eight widely recognized molecular properties, such as molecular mass and polar surface area. These two scores can be seen as more advanced validity metrics. More details are available in Appendix D.
Other performance metrics we use are the sample time (seconds) needed to generate a batch of $64$ molecules.
Results. The experimental results are presented in Tables 12, and 3, and they pertain to DMol, CDMol (Compressed DMol) proposed in Section 4, and CDiGress (Compressed DiGress). On QM9, DMol achieves the best performance compared to other baseline models. In contrast, the state-of-the-art discrete diffusion model, DiGress, shows a significant deficiency in novelty, indicating that most of the generated molecules are duplicates of those already present in the training set. CDMol retains the V and U scores of DMol, but significantly improves the N score. This is because ring compression makes it more likely for molecules to exhibit diverse structures.
On the MOSES and GUACAMOL datasets, our method achieves a performance comparable to that of the autoregressive model GraphInvent, which leverages chemical software and reinforcement learning for fine-tuning. DMol and CDMol also significantly outperform DiGress (note that the V scores of DiGress on MOSES and Guacamol were reported as $85.7$% and $85.2$%; however, we report the results of our own simulations, for which there is a discrepancy of roughly $1$% against DiGress - such discrepancy is not unexpected, since the output results are random). Although JT-VAE demonstrates excellent V performance, it involves a very complex reconstruction process and potentially suffers from issues associated with VAE-based generative models, such as coarse-graining (which manifests itself as fairly uniform scaffolds of the generated molecules). In contrast, DMol can directly sample molecular graphs with flexible control over the number of generated nodes.
Table 4 presents a comparison between our model and DiGress with respect to the CL score, QED, sample time and approximate training time. The results show that the average CLscore and QED of molecules generated by our model are slightly but uniformly higher than those of DiGress, and that the training and sampling speed are significantly faster. This advantage stems from the fact that our noise model adds noise to subgraphs rather than the entire graph, allowing the denoising model to converge more quickly. Also, our diffusion process uses a number of time steps proportional to the number of nodes in the training graph which further contributes to the reduced execution time. As a parting remark, we point out that although ring compression requires a (simple) decoding step, it significantly reduces the number of sampling steps needed since it effectively makes the compressed molecule graph smaller. However, when ring compression is combined with classical DiGress diffusion, since the total number of sampling steps is fixed, the decoding process actually increases the total sampling time. Therefore, compressed DMol also uniformly outperforms compressed DiGress. Additional results, pertaining to the GruM method (Jo et al., 2024), which were omitted in the main text due to space limitations, along with an ablation study of DMol, are available in Appendix E.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Ring Compression
Since each dataset has a different size of graph and complexity, we choose a different number of rings K to be converted to supernodes. We select $3$ rings for QM9 4, $15$ rings for Moses 5, and $15$ rings for Guacamol 6.
Figure 4: The $3$ selected carbon rings in QM9.
Figure 5: The $15$ selected carbon rings in Moses.
Figure 6: The $15$ selected carbon rings in Guacamol.
Appendix B “Converting” DMol Diffusion into Digress Diffusion
The following modifications can be applied to DMol to obtain Digress:
1. Fix the maximum number of sampling steps to a constant length instead of varying based on the number of nodes.
2. Scale the number of nodes to be changed at each step by a factor of $\sum_{i}p^{x}_{i}(1-p^{x}_{i})$.
3. Set the noising processes for edges and nodes to be independent, i.e., select edges from the entire graph instead of restricting to the subgraph formed by the selected nodes.
4. Scale the number of edges to be changed at each step by a factor of $\frac{\sum_{j}p^{e}_{j}(1-p^{e}_{j})}{r}$.
5. When generating random scores for nodes and edges to determine the objects to be selected, multiply the random scores for nodes by a correction weight of $1-p^{x}_{i}$, and multiply the random scores for edges by $1-p^{e}_{j}$.
These modifications are straightforward to implement in practice and can be achieved by simply adjusting model hyperparameters. However, Digress cannot be easily converted into DMol.
Appendix C Further comparisons between DMol and DiGress
C.1 Noise Distribution Evolution
In DiGress, the node type distribution remains fixed at all forward time steps because the initial node type distribution is the marginal distribution $\mathbf{m}_{X}$. This is easy to see because $p_{0}=\mathbf{m}_{X}$, so $p_{1}=\overline{\alpha}^{1}p_{0}+\overline{\beta}^{1}\mathbf{m}_{X}=\mathbf{m}_{X}$, and by induction, $p_{t}=\mathbf{m}_{X}$ for all $t$, where $p_{t}$ is the node type distribution at step $t$. In DMol, the node type distribution changes over time. For a node with type $i$ selected for perturbation, the probability of transitioning to type $j\neq i$ is $P(j|i_{\text{selected}})=\frac{p_{X}(j)}{1-p_{X}(i)}$. This leads to a different recurrence relation for the probability distribution over time.
C.2 Efficiency
To compare the efficiency of DMol and DiGress, we compare the time step increment required to change exactly one node under both settings. In DMol, the number of nodes changing at time $t$ is $N(t)=(1-\alpha^{t})\cdot n$. To find $\delta t_{DMol}$ such that $\delta N(t)=1$, we need
$N(t+\delta t)-N(t)=1$
The number of nodes that change at time $t+\delta t$ equals
$N(t+\delta t)=(1-\alpha(t+\delta t))\cdot n$
For DMol, we use $T=kn$ and the cosine schedule for $\alpha$,
$\alpha(t)=\cos^{2}\left(\frac{0.5\pi(t/T+s)}{1+s}\right).$
Taking the derivative with respect to $t$ and using a small angle approximation for small $\delta t$, we obtain:
$\frac{d\alpha(t)}{dt}\approx-\frac{\pi\sin\left(\frac{\pi(t/T+s)}{1+s}\right)\cos\left(\frac{\pi(t/T+s)}{1+s}\right)}{T(1+s)}=-\frac{\pi\sin\left(\frac{2\pi(t/T+s)}{1+s}\right)}{2T(1+s)}$
For small $\delta t$, we can approximate:
$\alpha(t+\delta t)-\alpha(t)\approx\frac{d\alpha(t)}{dt}\cdot\delta t$
Therefore,
$N(t+\delta t)-N(t)=(\alpha(t)-\alpha(t+\delta t))\cdot n\approx-\frac{d\alpha(t)}{dt}\cdot\delta t\cdot n$
Setting this equal to 1 and solving for $\delta t$, we can have the value of $\delta t_{DMol}$:
$\delta t_{DMol}=\frac{1}{-\frac{d\alpha(t)}{dt}\cdot n}=\frac{2T(1+s)}{\pi n\sin\left(\frac{2\pi(t/T+s)}{1+s}\right)}$
For DiGress, we use a similar approach to find $\delta t_{Digress}$. First, the expected number of nodes changed by DiGress at time $t$ equals:
$N_{Digress}(t)=n\cdot(1-\alpha(t))\cdot\sum_{i}p^{x}_{i}\cdot(1-p^{x}_{i})$
Following a similar derivation as for DMol, but with a constant $T$ instead of $T=kn$, we arrive at
$\delta t_{Digress}=\frac{1}{-\frac{d\alpha(t)}{dt}\cdot n\cdot\sum_{i}p^{x}_{i}\cdot(1-p^{x}_{i})}=\frac{2T(1+s)}{\pi n\sin\left(\frac{2\pi(t/T+s)}{1+s}\right)\cdot\sum_{i}p^{x}_{i}\cdot(1-p^{x}_{i})}$
Assume next, for fairness of comparison, that the maximum number of diffusion step used by DMol and Digress are the same. Then, the ratio of the minimum time steps of interest equals
$\frac{\delta t_{Digress}}{\delta t_{DMol}}=\frac{1}{\sum_{i}p^{x}_{i}\cdot(1-p^{x}_{i})}$
Because $\sum_{i}p^{x}_{i}\cdot(1-p^{x}_{i})<1$ for any probability distribution, we have that
$\frac{\delta t_{Digress}}{\delta t_{DMol}}>1$
This proves that $\delta t_{Digress}>\delta t_{DMol}$, or, in words, that DMol requires a smaller time step increment to change exactly one node. As a result, without restrictions on the diffusion times, the number of steps required by DMol to change a node will be smaller than that of required by Digress, which implies that DMol performs graph perturbations more efficiently.
C.3 Expressivity
Both DiGress and DMol start with the same underlying node and edge type distributions, but differ in how efficiently they explore the space of possible graphs. DMol tends to produce more diverse graphs within its reachable set because it makes larger localized changes. The Hamming distance between the initial graph $\mathcal{G}_{0}$ and the graph at time $t$, $\mathcal{G}_{t}$, is approximately $d_{H}(\mathcal{G}_{0},\mathcal{G}_{t})\sim\beta^{t}\cdot n$. This Hamming distance captures the difference between the number of node/edge types in the two graphs. For DiGress, the expected Hamming distance is $\mathbb{E}[d_{H}(\mathcal{G}_{0},\mathcal{G}_{t})]\sim\beta^{t}\cdot n\cdot\sum_{i\in\mathcal{C}_{X}}p_{X}(i)\cdot(1-p_{X}(i))$. Since $\sum_{i\in\mathcal{C}_{X}}p_{X}(i)\cdot(1-p_{X}(i))<1$, we conclude that DMol explores the graph space more efficiently per diffusion step.
Appendix D ChEMB Likeness and QED Scores
The ChEMBL-Likeness Score (CLscore) (Bühlmann & Reymond, 2020) is a metric used to assess the drug-likeness of a molecule. It is determined by analyzing how many of the molecule’s substructures are present in the ChEMBL dataset(Gaulton et al., 2016), a database of bioactive molecules with drug-like properties. Essentially, the CLscore quantifies the likelihood of a molecule being drug-like based on its substructure similarity to known compounds in ChEMBL.
Quantitative Estimation of Drug-likeness (QED) (Bickerton et al., 2012) is a computational metric used in drug discovery to assess how “drug-like” a compound is. It combines eight physicochemical properties, including molecular mass ($M_{r}$), octanol–water partition coefficient (ALOGP), number of hydrogen bond donors (HBDs), number of hydrogen bond acceptors (HBAs), molecular polar surface area (PSA), number of rotatable bonds (ROTBs), number of aromatic rings (AROMs), and number of structural alerts (ALERTS), into a single score ranging from 0 (least drug-like) to 1 (most drug-like). The score is measured through defined desirability functions of the eight properties. QED provides a balanced evaluation by weighting these properties based on their distributions in known drugs.
Appendix E Supplementary Experiments
E.1 General Graph Generation
We conducted experiments using the benchmark proposed by (Martinkus et al., 2022a), which comprises two datasets: SBM and Planar. Each dataset consists of 200 graphs. DMol’s ability to accurately model various properties of these graphs was assessed using metrics such as the degree distribution (Deg.), clustering coefficient distribution (Clus.), and orbit count distribution (Orb., the number of occurrences of substructures with 4 nodes), measured by the relative squared Maximum Mean Discrepancy (MMD).
MMD measures the discrepancy between two sets of distributions. The relative squared MMD is defined as follows:
$\frac{\text{MMD}^{2}\left(\mathcal{G}_{gen}\|\mathcal{G}_{test}\right)}{\text{MMD}^{2}\left(\mathcal{G}_{train}\|\mathcal{G}_{test}\right)},$
where $\mathcal{G}_{gen}$, $\mathcal{G}_{train}$, and $\mathcal{G}_{test}$ are the sets of generated graphs, training graphs, and test graphs.
In our experiments, we split the datasets into training, validation, and test sets with proportions of 64%, 16%, and 20%, respectively. The chosen hyperparameters were $k=2$ and $r=0.01$. The experimental results are presented in Table 5. These results demonstrate that DMol performs comparably to DiGress in generating general graphs.
Table 5: MMD Performance Comparison on SBM and Planar Graphs.
Dataset Model Deg ↓ Clus ↓ Orb ↓ Valid ↑
SBM GraphRNN 6.7 1.6 3.2 5.2
GRAN 14.3 1.7 2.0 25.5
SPECTRE 1.9 1.7 1.5 100.0
DiGress 1.7 1.6 1.7 66.7
DMol(Ours) 1.6 1.5 1.5 66.1
Planar GraphRNN 24.6 8.8 2534.0 0.0
GRAN 3.6 1.2 1.8 98.2
SPECTRE 2.6 2.5 2.5 100.0
DiGress 1.6 1.5 1.6 84.5
DMol(Ours) 1.8 1.6 1.4 83.9
E.2 Ablation Study
The ablation study highlights the performance improvements of DMol compared to DiGress due to the use of a new loss function that penalizes discrepancies in the counts of nodes and edges of training molecular graphs and sampled graphs; and performance improvements due to the use of special subgraph selection procedures. The results are summarized in Table 6. The model we used is DMol, and the dataset is GUACAMOL. In Table 6, “ONLY CE LOSS” stands for using only the cross-entropy loss (the loss used by DiGress), while “BOTH CE & MSE LOSS” refers to using both cross-entropy loss and mean squared error loss (the loss used by DMol). “SELECTING EDGES FROM THE WHOLE GRAPH” means selecting edges from the entire graph, whereas “SELECTING EDGES FROM THE INDUCED SUBGRAPH” means selecting edges only from the induced subgraph formed by the selected nodes. The results show that incorporating the MSE loss can improve the validity of the generated molecules. This is because the MSE loss penalizes discrepancies between the number of altered nodes and edges and the noise added during the forward process, resulting in more accurate predictions by the denoising model. Additionally, selecting edges from the induced subgraph significantly enhances the validity of the generated molecular graphs. This is achieved by making the sets of altered nodes and edges codependent. However, if edges are selected from the whole graph, the reduced number of diffusion steps significantly increases the learning difficulty for the denoising model, leading to lower validity in the generated molecular graphs.
Table 6: Ablation Study
Method V ↑ U ↑ N ↑
Only CE Loss 79.6 100 100
Both CE & MSE Loss 85.4 100 100
Selecting Edges from the Whole Graph 55.6 100 100
Selecting Edges from the Induced Subgraph 85.4 100 100
E.3 Selection of the parameter $k$
The hyperparameter $k$ determines the total number of diffusion steps since we set $T=kn$, where $n$ is the number of nodes. In practice, $k$ is typically set to $1$ or $2$ to reduce the number of diffusion steps and enhance performance (as explained in Figure 7). The latter observation can be intuitively explained as follows: the number of nodes modified at each time step increases over time, so that by the final time step, all nodes are changed. As $k$ increases, the x-axis of the diffusion process is effectively stretched by a factor of $k$, while the y-axis remains unchanged (Figure 7). Consequently, the number of diffusion steps required to modify a single node increases. This leads to a higher frequency of generated samples during the early stages of diffusion, when fewer nodes are altered; furthermore, fewer samples are generated during the “middle stages.” However, these additional early-stage samples introduce redundancy during the process of training the denoising network, ultimately degrading performance. Furthermore, these samples are generated at the cost of samples with a larger number of node changes, further degrading the performance.
Figure 7: Choosing a larger value of $k$ results in a less step-like (flatter) $N(t)$ function. As shown in the two figures, using $k=10$ leads to a more skewed distribution of $\delta t$ toward $N(t)$. This means that during sampling, the model is more likely to select $t$ during the early diffusion stages, which makes it harder for the neural network model to learn how to denoise.
In addition, to empirically evaluate the impact of $k$ on the performance of the diffusion model, we conducted experiments for different values of $k$. As shown in Table 7, the sample time is approximately proportional to $k$. Moreover, setting $k=2$ yields the highest validity.
[…]
1 85.1 100 100 2.53
2 85.6 100 100 5.28
3 84.9 100 100 7.81
4 84.7 100 100 10.37
5 84.4 100 100 12.69
10 83.4 100 100 24.76
Table 7: DMol with different values of $k$ on MOSES dataset
E.4 Comparison with GruM
We also compare CDMol to GruM (Jo et al., 2024) on the 2D molecule generation task. Since GruM was designed for small-scale datasets, we ran both methods on QM9 to ensure fairness. As shown in Table 8, GruM achieves slightly higher validity but significantly lower uniqueness and novelty scores. Moreover, GruM does not scale well, and we were unable to run it to completion on Guacamol using hardware that successfully supports both DiGress and DMol.
[…]
GruM 99.4 85.6 26.1 85.2 22.2
CDMol 98.3 97.7 78.2 96.0 75.1
Table 8: Comparison between CDMol and GruM on QM9 dataset
Appendix F Examples of Generated Molecules.
Figure 8 shows the sampling process.
Figure 9, 10, 11 show examples of molecule generation.
Figure 8: Sampling process.
Figure 9: Molecular graphs sampled from DMol trained on the QM9 dataset.
Figure 10: Molecular graphs sampled from DMol trained on the MOSES dataset.
Figure 11: Molecular graphs sampled from DMol trained on the GUACAMOL dataset.
Appendix G Docking Results
We used AutoDock Vina version 1.2.0 (Goodsell & Olson, 1990) to visualize the docking of two of the top-scoring generated molecules (with respect to the ChEMBL likeness) that also exhibit strong biding affinities on the 1PXH and 4MQS proteins when compared to reference drug ligants. The 4MQS complex represents the active human M2 muscarinic acetylcholine receptor bound to the antagonist iperoxo, while 1PXH represents the protein tyrosine phosphatase 1B with bidentate inhibitor compound 2. The free energy results are shown in Figures 12 and 13, and they indicate that the generated molecules can signficantly outperform or underperform existing reference drugs.
(a) Reference: Free energy = $-8.7$ kcal/mol
(b) Generated drug 1: Free energy = $-6.4$ kcal/mol
(c) Generated drug 2: Free energy = $-6.0$ kcal/mol
Figure 12: The reference molecule (a) and two of our generated molecules (b) and (c) docked at the protein receptor 1PXH, and the corresponding free energies. The generated molecules significantly underperform with respect to the reference due to their smaller sizes/masses. As a result, molecular weight constraints have to be considered as part of the design process.
(a) Reference: Free energy $-6.4$ kcal/mol
(b) Generated drug 1: Free energy = $-6.6$ kcal/mol
(c) Generated drug 2: Free energy = $-9.1$ kcal/mol
Figure 13: The reference molecule (a) and two of our generated molecules (b) and (c) docked at the protein receptor 4MQS, and the corresponding free energies. The generated molecules significantly outperform with respect to the reference, and the sizes/masses of the drug molecules are comparable. This further confirms the importance of focusing on the right molecular weight and size of the generated molecules.